# UniUIR: Considering Underwater Image Restoration as An All-in-One Learner

### v1

@@ Abstract
Existing underwater image restoration (UIR) methods generally only handle color distortion or jointly address color and haze issues, but they often overlook the more complex degradations that can occur in underwater scenes. To address this limitation, we propose a Universal Underwater Image Restoration method, termed as UniUIR, considering the complex scenario of real-world underwater mixed distortions as an all-in-one manner. To decouple degradation-specific issues and explore the inter-correlations among various degradations in UIR task, we designed the Mamba Mixture-of-Experts module. This module enables each expert to identify distinct types of degradation and collaboratively extract task-specific priors while maintaining global feature representation based on linear complexity. Building upon this foundation, to enhance degradation representation and address the task conflicts that arise when handling multiple types of degradation, we introduce the spatial-frequency prior generator. This module extracts degradation prior information in both spatial and frequency domains, and adaptively selects the most appropriate task-specific prompts based on image content, thereby improving the accuracy of image restoration. Finally, to more effectively address complex, region-dependent distortions in UIR task, we incorporate depth information derived from a large-scale pre-trained depth prediction model, thereby enabling the network to perceive and leverage depth variations across different image regions to handle localized degradation. Extensive experiments demonstrate that UniUIR can produce more attractive results across qualitative and quantitative comparisons, and shows strong generalization than state-of-the-art methods.
@@ I Introduction
Underwater images have proliferated across fields such as marine biology, underwater exploration, and underwater archaeology[1, 2]. They play a vital role in tasks like water body classification and ocean condition monitoring[3, 4], and in providing immersive visuals of underwater heritage and landscapes[5]. However, due to wavelength-dependent light attenuation and scattering by marine microorganisms[6], underwater scenes often exhibit multiple overlapping distortions, including haze, color shifts, low contrast, and blurred details. In addition, certain adverse imaging conditions, such as low-light environments, can further exacerbate the degradation of underwater image quality. These complex, mixed distortions pose significant challenges for both machine vision tasks and human perception, underscoring the need for effective underwater image restoration (UIR) task.
@@
Existing UIR methods can be broadly classified into three categories: visual prior-based, physical model-based, and deep learning-based approaches. Visual prior-based methods focus on adjusting image pixel values such as contrast, brightness, and saturation [8, 9], but fail to account for the physical degradation processes specific to underwater environments. Physical model-based methods [9, 10, 11] mainly estimate physical medium parameters of the underwater imaging process, yet their adaptability is limited by environmental complexity, resulting in reduced robustness. Recently, deep learning-based approaches [7, 12, 13, 14, 15, 16] have emerged, demonstrating notable improvements through end-to-end optimization. However, these methods often assume that all distortions can be handled simultaneously, typically focusing on specific types of distortions while neglecting the interactions and conflicts between them. As a result, their performance is limited in complex scenes with multiple simultaneous distortions. As illustrated in Fig. 2(a), many of these methods rely on physical modeling [17, 18] or jointly optimized depth estimation networks [14, 19, 20], which present several challenges: 1) Depth estimation networks require paired ground-truth depth data for optimization, which is difficult and costly to obtain. 2) Physical models rely on idealized assumptions, making them potentially inadequate for handling the complexity of real-world underwater environments, often fail to handle distortions such as detail blurring or low-light conditions. In contrast, our method (Fig. 2(b)) emphasizes representing degradation priors and leverages the capabilities of large-scale pre-trained depth prediction models. By incorporating an encoder-decoder framework equipped with a mixture-of-experts, our approach effectively addresses multiple distortions, including extreme low-light underwater scenarios, demonstrating robust and comprehensive restoration performance.
@@
In this study, we aim to systematically address the aforementioned challenges. Specifically, we propose a novel universal underwater image restoration method, termed UniUIR, designed to tackle the complex real-world scenario of mixed underwater distortions in an all-in-one manner. By analyzing the distortion characteristics and distributions in existing underwater image datasets, we develop MMoE-UIR, a Mamba Mixture-of-Experts-based restoration backbone that effectively captures and addresses diverse degradation patterns. To further enhance restoration performance, we design a spatial-frequency prior generator (SFPG) capable of extracting degradation priors from both spatial and frequency domains while adaptively selecting the most relevant task-specific prompts based on image content. Additionally, to address inconsistencies between distortions across different image regions, we incorporate scene depth information using the Depth Anything V2 [21] model, enabling the network to better distinguish and focus on visual discrepancies between foreground and background elements. Finally, to recover fine-grained details and mitigate information loss inherent in degradation priors, we leverage the powerful generative capabilities of diffusion models [22, 23]. The latent condition diffusion model is introduced to refine the restoration process, effectively modeling complex distributions and enabling the reconstruction of high-quality images with accurate and visually consistent details.
@@
❑ To leverage the interdependencies among various degradations in underwater environments, we developed the Mamba Mixture-of-Experts module. This module facilitates each expert in recognizing specific degradations and collectively inferring task-specific priors, while preserving both local and global feature representations.
@@ II-A Underwater Image Restoration
Recently, deep learning–based methods [7, 12, 13, 15, 26, 18, 19, 16, 27, 28] have shown significant potential in underwater image restoration (UIR). These methods leverage advanced neural network architectures and techniques to address challenges in image degradation. For instance, Waternet [7] integrates enhanced input modes with confidence maps to generate improved images, while Ucolor [12] employs multi-color space embedding guided by medium transmission to enhance detail and color restoration. Semi-UIR [26] combines contrastive and semi-supervised learning to improve performance and robustness, and HCLR-net [16] uses hybrid contrastive learning with locally randomized perturbations to enhance image quality. Despite these advancements, deep learning-based methods often assume that all distortions can be addressed simultaneously. This assumption leads to a focus on main distortion, for example, color cast or low contrast, while overlooking interactions and conflicts between multiple distortions. Consequently, their performance remains limited in complex underwater scenes with diverse and simultaneous degradations.
@@ II-B All-in-One Image Restoration
All-in-one image restoration [29, 30, 31, 32, 33, 34] has emerged as a promising approach in low-level vision tasks, aiming to recover clean images from multiple degraded inputs using a unified model. Compared to task-specific [35, 36, 37, 38, 39, 40] and general [41, 42, 43, 44, 45] restoration methods, all-in-one restoration offers advantages in model storage efficiency and practical applications. The main challenge lies in designing a single architecture capable of addressing diverse degradation types. To tackle this, methods like AirNet [29] employ contrastive learning for discriminative degradation representations, while IDR [30] uses a two-stage, ingredients-oriented approach. Vision prompts, as used in PromptIR [31] and ProRes [32], further enhance the handling of diverse degradations. More recently, MPerceiver [34] has leveraged pre-trained large-scale vision models to excel in all-in-one restoration tasks.
@@ II-C Diffusion Models for Image Restoration
Diffusion models (DM) [22, 23], initially developed for generative tasks, have demonstrated strong potential in image restoration by addressing inverse problems through progressive denoising. Their capability to model complex data distributions has led to state-of-the-art results in low-level vision tasks such as super-resolution [46, 47] and low-light image enhancement [48, 49]. Recently, the generative power of DM has attracted attention in underwater image restoration (UIR) [50, 51]. For example, Tang et al. [50] proposed a diffusion-based underwater image enhancement method, while UIEDP [51] introduced a framework incorporating diffusion priors, formulating UIR as posterior sampling conditioned on degraded inputs.
Although DM excel at generating high-quality images, they face notable challenges in UIR, including low efficiency, limited adaptability to diverse degradation types, insufficient integration of environmental priors, and constrained generalization capabilities. To address these limitations, our UniUIR incorporates a latent condition diffusion model to refine the restoration process and reconstruct high-quality images with consistent and precise details, offering a more effective solution for UIR.
@@ III-A Overview Pipeline
$\mathcal{L}_{\text{stageII}}=\mathcal{L}_{{1}}(\mathbf{X}_{gt},\mathbf{X}_{hq})+\mathcal{L}_{\text{diff}},\enspace\mathcal{L}_{\text{diff}}=\mathcal{L}_{{1}}(\mathbf{Z},\mathbf{\hat{Z}}).$ (4)
@@ III-B MMoE-UIR
$\mathbf{\hat{F}}=\mathcal{M}\Big{(}\text{LN}(\mathbf{F}_{d})\odot\text{Softmax}\big{(}\mathcal{C}_{3\times 3}(\mathbf{D}_{lq}^{n})\big{)}\Big{)}+\mathbf{F}_{d},$ (6)
where $\mathbf{\hat{F}}$ represents the output of the MMoEB; $\mathcal{M}(\cdot)$ denotes the Water Mixture-of-Experts operation; $\mathcal{C}_{3\times 3}$ denotes 3$\times$3 convolution, followed by a ReLU function; $\mathbf{D}_{lq}$ is generated from the pre-trained Depth Anything V2 model. Each MMoEM contains multiple MMoEBs.
@@ III-C Spatial-Frequency Prior Generator
Reconstructing image details in severely degraded regions is challenging due to the significant loss of high-frequency information caused by complex distortions. To restore these missing details, external guidance from high-quality underwater images is necessary. To achieve this, we use a Spatial-Frequency Prior Generator (SFPG) to extract a compact prior $\mathbf{Z}$ from $\mathbf{X}_{lq}$ and $\mathbf{X}_{gt}$, injecting it into MMoE-UIR to provide correct information for restoring the lost details. Specifically, as shown in Fig. 3, SFPG first employs PixelUnshuffle and 3$\times$3 convolutions with LReLU activation functions to downsample the concatenated images $\mathbf{X}_{c}\in\mathbb{R}^{H\times W\times 6}$ and capture effective information, resulting in $\mathbf{X}_{s}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C^{\prime}}$. Subsequently, two parallel branches of three residual blocks with 3$\times$3 convolutions and LReLU activation functions are utilized to extract features from the amplitude and phase of $\mathbf{X}_{s}$.
$\mathbf{X}_{\mathcal{A}},\mathbf{X}_{\mathcal{P}}=\mathcal{F}(\mathbf{X}_{s}),\enspace\tilde{\mathbf{X}}_{\mathcal{A}},\tilde{\mathbf{X}}_{\mathcal{P}}=\text{ResBlock}(\mathbf{X}_{\mathcal{A}},\mathbf{X}_{\mathcal{P}}),$ (9)
@@
2:  Parameters: $N$ Experts $\mathcal{E}$, Router $\mathcal{G}$, and top-k expert
@@ III-D Latent Conditional Diffusion Model
Set Training UIEB-T[7] Paired 800 Wreck, People, Fish, Reef, and Coral (259$\times$194), $\cdots$, (2180$\times$1447)
@@
U45[52] Unpaired 45 Wreck, People, Fish, Reef, and Coral 256$\times$256
SQUID-16[53] Unpaired 16 Wreck and Reef 512$\times$512
Challenge-60[7] Unpaired 60 People, Fish, Reef, Medusa (259$\times$194), $\cdots$, (2000$\times$1124)
UCCS[54] Unpaired 300 Reef and Sea Urchin 400$\times$300
EUVP-330[55] Unpaired 330 People, Fish, Reef, and Coral (320$\times$240), $\cdots$, (960$\times$540)
Set Training LSUI-T[13] Paired 3879 Fish, Reef, Coral, and Sea Urchin (256$\times$256), $\cdots$, (1280$\times$1024)
@@ IV-A Experimental Setup
Evaluation Metrics. We evaluate performance using reference metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [56], as well as non-reference metrics such as the Underwater Colour Image Quality Evaluation Metric (UCIQE) [57], Underwater Image Quality Measure (UIQM) [58], and URanker [15]. For the metrics PSNR, SSIM, UCIQE, UIQM, and URanker, higher scores indicate better performance. Conversely, for the LPIPS metric, lower scores are preferable.
Baselines. We compare our method against existing underwater image restoration methods, encompassing both traditional and deep learning approaches. These include a traditional method MLLE [25] and the following deep learning models: WaterNet [7], FUnIE [59], S-UWNet [60], Ucolor [12], PUIE-Net [61], Ushape [13], PUGAN [14], NU2Net [15], Semi-UIR [26], GUPDM[18], HCLR-Net [16]. Additionally, we compared our approach with the all-in-one image restoration method PromptIR[31] and the MambaIR[43] based on the Mamba architecture [62].
Implementation Details. All experiments were conducted using PyTorch on 8 NVIDIA RTX 3090 GPUs and optimized using the AdamW optimizer with momentum parameters (0.9, 0.999). During training stage I and II, UniUIR was trained for 50K and 200K iterations, respectively, with an initial learning rate of $2\times 10^{-4}$, gradually reduced to $1\times 10^{-6}$ using cosine annealing. The batch size was set to 8. MMoE-UIR utilized a 4-level encoder-decoder architecture with [3, 5, 6, 6] MMoEBs and channel dimensions of [32, 64, 128, 256] for levels 1 through 4. The channel numbers ${\hat{C}}$ in both SFPG and SFPG* were set as 256. A linear noise schedule was employed for diffusion process with parameters $\alpha_{1}=0.99$, $\alpha_{T}=0.1$, and a time step of $T=4$. Depth information was extracted using the large version of the Depth Anything V2 model [21]. For fair comparisons, all deep learning-based UIR methods were retrained using the original settings specified in their respective papers. Input images were cropped to 128$\times$128 pixels, and random flipping was applied for data augmentation. The hyperparameters were set as $\lambda_{1}=0.1$ and $\lambda_{2}=0.5$.
@@
Figure 6: Visual comparison of restoration results on the non-reference datasets: U45 [52], SQUID-16 [53], Challenge-60 [7], UCCS [54], and EUVP-330 [55].

### v2

@@ Abstract
Existing underwater image restoration (UIR) methods generally only handle color distortion or jointly address color and haze issues, but they often overlook the more complex degradations that can occur in underwater scenes. To address this limitation, we propose a Universal Underwater Image Restoration method, termed as UniUIR, considering the complex scenario of real-world underwater mixed distortions as an all-in-one manner. To disentangle degradation-specific effects and capture their inter-correlations, we propose the Mamba Mixture-of-Experts module (MMoEM). Each expert specializes in distinct aspects of degradation, while gating mechanism dynamically routes features to appropriate experts. This design enables collaborative prior extraction and preserves global context, all within linear computational complexity. Building upon this foundation, to enhance degradation representation and address the task conflicts that arise when handling multiple types of degradation, we introduce the spatial-frequency prior generator. This module extracts degradation prior information in both spatial and frequency domains, and adaptively selects the most appropriate task-specific prompts based on image content, thereby improving the accuracy of image restoration. Finally, to more effectively address complex, region-dependent distortions in UIR task, we incorporate depth information derived from a large-scale pre-trained depth prediction model, thereby enabling the network to perceive and leverage depth variations across different image regions to handle localized degradation. Extensive experiments demonstrate that UniUIR can produce more attractive results across qualitative and quantitative comparisons, and shows strong generalization than state-of-the-art methods. Project page at UniUIR.
@@ I Introduction
Underwater images have proliferated across fields such as marine biology, underwater exploration, and underwater archaeology[1]. They play a vital role in tasks like water body classification and ocean condition monitoring[2], and in providing immersive visuals of underwater heritage and landscapes. However, due to wavelength-dependent light attenuation and scattering by marine microorganisms[3], underwater scenes often exhibit multiple overlapping distortions, including haze, color shifts, low contrast, and blurred details. In addition, certain adverse imaging conditions, such as low-light environments, can further exacerbate the degradation of underwater image quality. These complex, mixed distortions pose significant challenges for both machine vision tasks and human perception, underscoring the need for effective underwater image restoration (UIR) task.
@@
Existing UIR methods can be broadly classified into three categories: visual prior-based, physical model-based, and deep learning-based approaches. Visual prior-based methods focus on adjusting image pixel values such as contrast, brightness, and saturation [5], but fail to account for the physical degradation processes specific to underwater environments. Physical model-based methods [6, 7] mainly estimate physical medium parameters of the underwater imaging process, yet their adaptability is limited by environmental complexity, resulting in reduced robustness. Recently, deep learning-based approaches [8, 9, 10, 11, 12] have emerged, demonstrating notable improvements through end-to-end optimization. However, these methods often assume that all distortions can be handled simultaneously, typically focusing on specific types of distortions while neglecting the interactions and conflicts between them. As a result, their performance is limited in complex scenes with multiple simultaneous distortions. As illustrated in Fig. 2(a), many of these methods rely on physical modeling [7, 13] or jointly optimized depth estimation networks [10, 14], which present several challenges: 1) Depth estimation networks require paired ground-truth depth data for optimization, which is difficult and costly to obtain. 2) Physical models rely on idealized assumptions, making them potentially inadequate for handling the complexity of real-world underwater environments, often fail to handle distortions such as detail blurring or low-light conditions. In contrast, our method (Fig. 2(b)) emphasizes representing degradation priors and leverages the capabilities of large-scale pre-trained depth prediction models. By incorporating an encoder-decoder framework equipped with a mixture-of-experts, our approach effectively addresses multiple distortions, including extreme low-light underwater scenarios, demonstrating robust and comprehensive restoration performance.
@@
In this paper, we aim to systematically address the aforementioned challenges. Specifically, we propose a novel universal underwater image restoration method, termed UniUIR, designed to tackle the complex real-world scenario of mixed underwater distortions in an all-in-one manner. By analyzing the distortion characteristics and distributions in existing underwater image datasets, we develop MMoE-UIR, a Mamba Mixture-of-Experts-based restoration backbone that effectively captures and addresses diverse degradation patterns. To further enhance restoration performance, we design a spatial-frequency prior generator (SFPG) capable of extracting degradation priors from both spatial and frequency domains while adaptively selecting the most relevant task-specific prompts based on image content. Additionally, to address inconsistencies between distortions across different image regions, we incorporate scene depth information using the Depth Anything V2 [15] model, enabling the network to better distinguish and focus on visual discrepancies between foreground and background elements. Finally, to recover fine-grained details and mitigate information loss inherent in degradation priors, we leverage the powerful generative capabilities of diffusion models [16, 17]. The latent condition diffusion model is introduced to refine the restoration process, effectively modeling complex distributions and enabling the reconstruction of high-quality images with accurate and visually consistent details.
@@
❑ To leverage interdependencies among underwater degradations, we developed a Mamba Mixture-of-Experts module, which dynamically routes features to experts specializing in different degradation aspects and fuses their outputs to infer task-specific priors.
@@ II-A Underwater Image Restoration
Recently, deep learning based methods [4, 8, 9, 11, 20, 13, 14, 12, 21, 22] have shown significant potential in underwater image restoration (UIR). These methods leverage advanced neural network architectures and techniques to address challenges in image degradation. For instance, Waternet [4] integrates enhanced input modes with confidence maps to generate improved images, while Ucolor [8] employs multi-color space embedding guided by medium transmission to enhance detail and color restoration. Semi-UIR [20] combines contrastive and semi-supervised learning to improve performance and robustness, and HCLR-net [12] uses hybrid contrastive learning with locally randomized perturbations to enhance image quality. Despite these advancements, deep learning-based methods often assume that all distortions can be addressed simultaneously. This assumption leads to a focus on main distortion, such as color cast or low contrast, while overlooking interactions and conflicts between multiple distortions. Consequently, their performance remains limited in complex underwater scenes with diverse and simultaneous degradations.
@@ II-B All-in-One Image Restoration
All-in-one image restoration [23, 24, 25, 26, 27] has emerged as a promising approach in low-level vision tasks, aiming to recover clean images from multiple degraded inputs using a unified model. Compared to task-specific [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] and general [39, 40, 41, 42] restoration methods, all-in-one restoration offers advantages in model storage efficiency and practical applications. The main challenge lies in designing a single architecture capable of addressing diverse degradation types. To tackle this, methods like AirNet [23] employ contrastive learning for discriminative degradation representations. Vision prompts, as used in PromptIR [24] and ProRes [25], further enhance the handling of diverse degradations. More recently, MPerceiver [27] and Perceive-IR[26] have leveraged pre-trained large-scale vision models to excel in all-in-one restoration tasks.
@@ II-C Diffusion Models for Image Restoration
Diffusion models (DMs) [16, 17], originally developed for generative tasks, have demonstrated strong potential in image restoration by addressing inverse problems through progressive denoising. Their capability to model complex data distributions has led to state-of-the-art results in low-level vision tasks such as super-resolution [43, 44]. Recently, the generative power of DM has attracted attention in underwater image restoration (UIR) [45, 46]. For example, Tang et al. [45] proposed a diffusion-based underwater image enhancement method, while UIEDP [46] introduced a framework incorporating diffusion priors, formulating UIR as posterior sampling conditioned on degraded inputs.
While DMs offer a principled way to model uncertainty, they still face practical limitations in UIR, including low sampling efficiency, poor adaptability to diverse degradation types, insufficient integration of environmental priors (e.g., depth, scattering properties), and limited generalization. To address these issues, we propose UniUIR, a unified framework built upon a novel Latent Conditional Diffusion Model (LCDM). Unlike standard DMs, LCDM explicitly incorporates high-level semantic and structural priors into the denoising process, enabling degradation-aware, fine-grained restoration. By conditioning on these priors in the latent space, our model achieves enhanced perceptual quality and structural coherence, while improving efficiency and generalization through targeted guidance.
@@ III-A Overview Pipeline
[…]
@@ III-B MMoE-UIR
$\mathbf{\hat{F}}_{d}=\text{LN}(\mathbf{F}_{d})\odot\text{Softmax}\big(\mathcal{C}_{3\times 3}(\mathbf{D}_{lq}^{n})\big),$ (6)
$\mathbf{\hat{F}}=\mathcal{M}\big(\mathbf{\hat{F}}_{d}\odot\text{Linear}(\mathbf{Z}_{n})+\text{Linear}(\mathbf{Z}_{n})\big)+\mathbf{F}_{d},$ (7)
where $\mathbf{\hat{F}}$ represents the output of the MMoEB; $\mathcal{M}(\cdot)$ denotes the Water Mixture-of-Experts operation; $\mathcal{C}_{3\times 3}$ denotes $\times3$ convolution, followed by a ReLU function; $\mathbf{D}_{lq}$ is generated from the pre-trained Depth Anything V2 model. Each MMoEM contains multiple MMoEBs.
@@ III-C Spatial-Frequency Prior Generator
Reconstructing image details in severely degraded regions is challenging due to the significant loss of high-frequency information caused by complex distortions. To restore these missing details, external guidance from high-quality underwater images is necessary. To achieve this, we use a Spatial-Frequency Prior Generator (SFPG) to extract a compact prior $\mathbf{Z}$ from $\mathbf{X}_{lq}$ and $\mathbf{X}_{gt}$, injecting it into MMoE-UIR to provide correct information for restoring the lost details. Specifically, as shown in Fig. 3, SFPG first employs PixelUnshuffle and $\times3$ convolutions with LReLU activation functions to downsample the concatenated images $\mathbf{X}_{c}\in\mathbb{R}^{H\times W\times 6}$ and capture effective information, resulting in $\mathbf{X}_{s}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C^{\prime}}$. Subsequently, two parallel branches of three residual blocks with $\times3$ convolutions and LReLU activation functions are utilized to extract features from the amplitude and phase of $\mathbf{X}_{s}$.
$\mathbf{X}_{\mathcal{A}},\mathbf{X}_{\mathcal{P}}=\mathcal{F}(\mathbf{X}_{s}),\kern 5.0pt\tilde{\mathbf{X}}_{\mathcal{A}},\tilde{\mathbf{X}}_{\mathcal{P}}=\text{ResBlock}(\mathbf{X}_{\mathcal{A}},\mathbf{X}_{\mathcal{P}}),$ (10)
@@
2:Parameters: NN Experts $\mathcal{E}$, Router $\mathcal{G}$, and top-k expert
@@ III-D Latent Conditional Diffusion Model
Set
@@
[…]
U45[47] Unpaired 45 Wreck, People, Fish, Reef, and Coral $\times256$
SQUID-16[48] Unpaired 16 Wreck and Reef $\times512$
[…]
Set
[…]
@@ IV-A Experimental Setup
Evaluation Metrics. We evaluate performance using reference metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [51], as well as non-reference metrics such as the Underwater Colour Image Quality Evaluation Metric (UCIQE) [52], Underwater Image Quality Measure (UIQM) [53], URanker [11], Underwater Image Fidelity (UIF) [54], and Underwater Image Quality Index (UIQI) [55]. For the metrics PSNR, SSIM, UCIQE, UIQM, URanker, UIF, and UIQI, higher scores indicate better performance. Conversely, for the LPIPS metric, lower scores are preferable.
Baselines. We compare our method against existing underwater image restoration methods, encompassing both traditional and deep learning approaches. These include a traditional method MLLE [19] and the following deep learning models: WaterNet [4], FUnIE [56], S-UWNet [57], Ucolor [8], PUIE-Net [58], Ushape [9], PUGAN [10], NU2Net [11], Semi-UIR [20], GUPDM[13], and HCLR-Net [12]. Additionally, we compared our approach with the all-in-one image restoration method PromptIR[24] and the MambaIR[40] based on the Mamba architecture [59].
Implementation Details. All experiments were conducted using PyTorch on 8 NVIDIA RTX 3090 GPUs and optimized using the AdamW optimizer with momentum parameters (0.9, 0.999). During training stage I and II, UniUIR was trained for 50K and 200K iterations, respectively, with an initial learning rate of $2\times 10^{-4}$, gradually reduced to $1\times 10^{-6}$ using cosine annealing. The batch size was set to 8. MMoE-UIR utilized a 4-level encoder-decoder architecture with [3, 5, 6, 6] MMoEBs and channel dimensions of [32, 64, 128, 256] for levels 1 through 4. The channel number $^{\hat{C}}$ for both SFPG and SFPG* was 256. Task-related prompts (256-dimensional) were initialized as learnable parameters. Within the W-MoE module, we employed 3 experts and activated the top-$k$ (kk=2) for each input. A linear noise schedule was employed for diffusion process with parameters $\alpha_{1}=0.99$, $\alpha_{T}=0.1$, and a time step of $T=4$. Depth information was extracted using the large version of the Depth Anything V2 model [15]. For fair comparisons, all learning-based UIR methods were retrained using the original settings specified in their respective papers. Input images were cropped to $\times128$ pixels, and random flipping was applied for data augmentation. The hyperparameters were set as $\lambda_{1}=0.1$ and $\lambda_{2}=0.5$.
@@
Figure 6: Visual comparison of restoration results on the non-reference underwater image datasets: U45 [47], SQUID-16 [48], Challenge-60 [4], UCCS [49], and EUVP-330 [50].
Figure 7: Comparison of UIF [54] and UIQI [55] scores for state-of-the-art methods on non-reference underwater image datasets.